OpenStudy (anonymous):

I am lost here. Mulitply & simplify 64c^2/4c^2-24c+36*4c-12/8c

OpenStudy (anonymous):

To clarify, is this the expression? \[\frac{64c^2}{4c^2}-24c+36\times 4c - \frac{12}{8c}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Let's take the first part: \[\frac{64c^2}{4c^2}\]What do you get when you divide 64 by 4? What do you get when you divide \(c^2\) by \(c^2\)?

OpenStudy (anonymous):

14c?

OpenStudy (anonymous):

Take each question individually, perhaps it will be easier.

OpenStudy (anonymous):

Im still lost